OpenStudy (anonymous):

1.) square root of x+2=square root of 3x=8 2.) cube root of 2x+5=-3 3.) cube root of 3x+1=5 help.. :D

OpenStudy (lalaly):

is tht 3x-8

OpenStudy (anonymous):

yes.. :)

OpenStudy (lalaly):

1) square both sides x+2=3x-8 x-3x=-8-2 -2x=-10 x=5

OpenStudy (lalaly):

to get rid of a square root u square both sides of the equation same goes to the cubic root but u cube both sides,,, try doing the next one

OpenStudy (anonymous):

tnx! :D

OpenStudy (lalaly):

ur wwelcme:)